Find MPC, when multiplier (K) = 1.


1
Expert's answer
2020-07-22T08:37:33-0400

Answer


"MPC = 0"


Workings

MPC stands for marginal propensity to consume. It is the change is consumer spending brought out by a change in income.

"MPC = \dfrac {∆C} {∆Yd}" Where ∆C represents a change in consumption and ∆Yd is a change in disposable income.


Multiplier (k) represents a change in income due to a change in injections in the circular flow of income.


MPC is used to derive the multiplier (k) as follows:

"K = \dfrac {1} {MPS}" Where MPS is the marginal propensity to save.


"MPC + MPS = 1"


Hence,

"MPS = 1 - MPC"


Therefore,

"K = \dfrac {1} {1 - MPC}"


It is given that K = 1, and hence,

"1 = \dfrac {1} {1 - MPC}"


Cross multiplying gives:

"1(1 - MPC) = 1"

"=> 1 - MPC = 1"

"=> 1 - 1 = MPC"

"=> MPC = 0"


So, when k = 1, MPC = 0